OpenStudy (taylor):

x-2y=8 can i get help on trying to find the answer

OpenStudy (anonymous):

i love this

OpenStudy (anonymous):

what are you solving for?

OpenStudy (taylor):

can you help me out i kinda suck at math

OpenStudy (anonymous):

x and y

OpenStudy (taylor):

im sovling for y

OpenStudy (anonymous):

First, change it to slope-intercept form: x-2y=8 (subtract x from both sides) -2y=8-x (divide both sides by -2) y=-4+.5x (rearrange the right side) y=.5x-4 Your .5 x, or 1/2, is your slope. It's rise over run. The -4 is your y-intercept. So, go to the point (0,-4). From there, go up one, over to the right two, up one, over to the right two. That's how you graph the equation. =)

OpenStudy (anonymous):

x-2y=8 -2y=8-x 2y=-8+x y=-4+1/2x y=1/2x-4

OpenStudy (anonymous):

thom WHT?? is That

OpenStudy (anonymous):

thom WHT?? is That

OpenStudy (anonymous):

is that what you r doing

OpenStudy (taylor):

i thinks so let me see

OpenStudy (anonymous):

ok if not i will try agine

OpenStudy (taylor):

on my paper it just says x-2y=8 im not graphing anything and im solving for y

OpenStudy (anonymous):

ok i am srry

OpenStudy (anonymous):

i will try agine

OpenStudy (taylor):

ok

OpenStudy (anonymous):

ok her we go Do as follows: 2x + 2y = 8 Notice that you have a common factor of 2, therefore divide everything by 2 2x/2 + 2y/2 = 8/2 Which then becomes: x + y = 4 Now substract x from both sides to isolate the y: x + y - x = 4 - x x - x + y = 4 - x; collecting like terms 0 + y = 4 - x y = 4 - x REMEMBER: That when you are trying to remove a term from one side do the opposite on BOTH sides. Here I eliminated the x by substracting x on both sides.